OpenStudy (anonymous):

Good Morning. Can someone help me verify my answers for 2 problems. y^2+12y=-3(this a completeing the square problem.

OpenStudy (anonymous):

\[(y+6)^2=-3+36=33\] \[y+6=\sqrt{33}\] \[y=-6\pm\sqrt{33}\]

OpenStudy (anonymous):

great I that's what I got. thanks.